OpenStudy (anonymous):

Express as a polynomial \[(3x+2y)^2(3x-2y)^2\]

OpenStudy (anonymous):

(3x+2y)(3x+2y)(3x-2y)(3x-2y) simplify.

OpenStudy (anonymous):

Okay, I think I have an idea of what to do I'll let you know :=)

OpenStudy (shane_b):

If it helps, I believe the easiest way to simplify it would be to do it in 2 parts...so do (3x+2)(3x+2) then the other part...and the multiply them together.

OpenStudy (anonymous):

Okay... I took a totally different route... I didn't multiply the whole thing out if that's what your implying @Shane_B But here's what I did:\[=(3x+2y)(3x+2y)(3x-2y)(3x-2y)\]\[=(3x+2y)(3x-2y)(3x+2y)(3x-2y)\]And since that is the difference of two squares, then this becomes:\[=(9x^2-4y^2)(9x^2-4y^2)\]\[=(9x^2-4y^2)^2\]\[=81x^4-72x^2y^2+16y^4\]

OpenStudy (anonymous):

But thanks so much for your help @monokerous and @Shane_B

OpenStudy (anonymous):

yw.